# Mapping a dissipative quantum spin chain onto a generalized Coulomb gas

### v2

@@ 1 Introduction
The focus of this paper is on dissipative 1D quantum systems. It turns our that introducing dissipation amounts to replacing the previously mentioned quenched disorder with dynamical (annealed) disorder. Following the formalism developed by Caldeira, Leggett et al. [21, 22, 23, 24, 25], dissipation is introduced by coupling to a bath of phonons that is simple enough to be traced out, thus yielding an effective description of the system of interest. Although this type of bath was first studied by putting it in contact with a single degree of freedom such as a spin in the spin-boson model [26, 27], several many-body 1D systems were later proposed, ranging from spin chains [28, 29, 30, 31, 32, 33] and 1D bosons [34], to glassy systems [35, 36], and with various system-bath couplings [37, 38]. In this work, we concentrate on the well-known XXZ spin chain at zero magnetization and temperature. Each spin of the chain is linked to a phonon bath à la Caldeira and Leggett, which can be seen as a many-body extension of the spin-boson model. The baths are furthermore spatially correlated by bath-bath nearest neighbor interactions. To the best of the author’s knowledge, it is the first time that such correlated dissipation is studied, previous studies having concentrated on local independent baths [31, 32, 33] or a single global bath [39]. For a spin chain at zero magnetization and with uncorrelated local baths, it has been proposed in [33] that the system presents a BKT-like transition which is driven by dissipation for subohmic baths. In this article, we broaden the scope of this study by considering correlated baths. In order to fully understand the nature of the transition in our system, we follow an approach based on a mapping to a generalized Coulomb gas. Within this gas picture, we show that for superohmic baths, the transition is the exact BKT transition while, for subohmic baths, dissipation induces a BKT-like transition.
The manuscript is organized as follows. Section 2 gives a concise summary of the main results which are derived in the following sections. Section 3 presents the microscopic model and shows how its low-energy behavior can be described by a dissipative sine-Gordon field theory. In section 4, this field theory is mapped onto a generalized 2D Coulomb gas and its basic features are described. The core of the article, identifying the BKT-like transition, is done in section 5. Finally, a brief discussion of the results and concluding remarks are made in section 6. Some technical details have been relegated in appendices A to E. Throughout this article, we set $\hbar=k_{B}=1$.
@@ 3 Model
where $s>0$ is the bath exponent, $\alpha$ measures the strength of the coupling to the bath, and $1/\tau_{c}$ is a frequency cutoff above which $J(\Omega)=0$. The exact numerical prefactor in Eq. (5) is unimportant and can be found in appendix A. This form of $J(\Omega)$ accounts for many types of environments, ranging from acoustic-phonon baths to Kondo physics [26]. Following the classification from Caldeira and Leggett, we call $s=1$ an ohmic bath, while $10<s<$ 1 is a subohmic bath and $s>1$ is a superohmic bath. The rest of this section will be dedicated to deriving an effective action for the system. The procedure is similar to that outlined in [33, 31, 32] and relies on two key steps: first bosonizing the spin chain, and then integrating out the bath degrees of freedom.
@@ 3.2 Integrating out the bath degrees of freedom
where $K_{0}$ is a modified Bessel function of the second kind, and we have used the definition of the spectral function given in Eq. (4). Upon using the low-energy behavior of $J(\Omega)$ given in Eq. (5), one can show that $\mathcal{K}(x,\tau)\simeq\alpha\tau_{c}^{s-1}\left(\tau^{2}+x^{2}/v^{2}\right)^{-1-s/2}/u$ for $\sqrt{\tau^{2}+x^{2}/v^{2}}\gg\tau_{c}$ (see appendix A). Putting everything together, one arrives at the following effective interaction
@@
where $\mathcal{D}(x,\tau)=\left[(\tau/\tau_{c})^{2}+(u/v)^{2}(x/a)^{2}\right]^{-1-s/2}$ is the dimensionless dissipative kernel. This effective action is that of a Luttinger liquid (LL) with two types of interactions: a local interaction describing the spin chain’s internal dynamics and controlled by $g$, and a long-range bath-induced interaction controlled by $\alpha$. This dissipative interaction can be thought of as the bath storing information about a spin at position $x$ and time $\tau$, propagating it over a distance $x-x^{\prime}$ during a period $\tau-\tau^{\prime}$, and giving it back at time $\tau^{\prime}$ to the spin at position $x^{\prime}$. Notice that this very-strong non-Markovian effect was obtained through the exact integration of the bath and is a feature which cannot be reproduced by usual open quantum system techniques such as the Lindblad master equation. As mentioned previously, the effective action written in Eq. (3.2) has already been studied in the $v=0$ case [33] which corresponds to the limit of uncorrelated baths. In this case, the last term in Eq. (3.2) gets modified by changing the kernel into $\mathcal{D}(x,\tau)\propto\delta(x)/(\tau/\tau_{c})^{1+s}$  [^2] This limit cannot be obtained by taking $v\to 0$ in the kernel $\mathcal{D}(x,\tau)$, but one rather has to redo the computation from the beginning with $v=0$. This is because the Green’s function of the bath operator $-v^{2}\partial_{x}^{2}-\partial_{\tau}^{2}+\Omega_{\gamma}^{2}$ does not continuously deform to that of the local bath operator $-\partial_{\tau}^{2}+\Omega_{\gamma}^{2}$ when $v\to 0$.. On the other hand, the case of $v\gg 1$ is new and corresponds to the limit of Markovian, but spatially correlated, baths. In this limit, the bath-induced kernel becomes $\mathcal{D}(x,\tau)\propto\delta(\tau)/(x/a)^{1+s}$. These two limits for $v$ formally yield identical models up to an exchange of space and time (and a redefinition of $\alpha$ to absorb extra factors of $v$). Another remarkable model is acheived when $v=u$ which corresponds to an isotropic system enjoying an $O(2)$ symmetry in $(x,u\tau)$ space. For the majority of this paper we will mainly deal with finite $v$ and postpone the discussion of the limiting cases ($v$ small or large) to section 6.
@@ 6 Discussion and conclusion
The procedure followed in this article can be straightforwardly extended to other one-dimensional quantum systems described by a sine-Gordon-like effective action. Studies of an XXZ spin chain, at zero [33] and finite magnetization [31, 32], subject to local baths (i.e. $v=0$ in our model) can be readily understood in the gas picture. Adapting our analysis, one would find pairs of bath-induced particles with an orientation locked in the imaginary-time direction. While pairs of particles with fixed orientation might seem to be an artifact of this model, it is actually a feature shared by all field theories endowed with the statistical tilt symmetry (STS) or space-gauged shift symmetry, i.e. the invariance of the interacting part of the action $S-S_{\rm LL}$ under $\phi(x,\tau)\to\phi(x,\tau)+f(x)$. In the gas picture, the STS indeed requires that the total charge along any $x={\rm cst}$ slice of the gas must be zero, thus strongly restricting the possible relative orientations of particles. This result is easily shown by considering the Boltzmann weight $\mathcal{W}\propto\langle\prod_{j}\exp[i4q_{j}\phi(r_{j})]\rangle_{\rm LL}$ associated with a configuration of particles of charges $q_{j}$ at positions $r_{j}$. Using the STS, one gets that $\mathcal{W}\propto\exp[i4\sum_{j}q_{j}f(x_{j})]\langle\prod_{j}\exp[i4q_{j}\phi(r_{j})]\rangle_{\rm LL}$ for all functions $f(x)$. This implies that allowed particle configurations must be neutral along any $x={\rm cst}$ slice of the 2D space. The STS appears generically in the replicated action of one-dimensional systems with quenched disorder, as in the well-known Giamarchi-Schulz model [16, 19].
@@ Acknowledgements
Appendix A Computation of the bath kernel
@@
[…]

### v3

@@ 1 Introduction
The focus of this paper is on dissipative 1D quantum systems. It turns our that introducing dissipation amounts to replacing the previously mentioned quenched disorder with dynamical (annealed) disorder. Following the formalism developed by Caldeira, Leggett et al. [21, 22, 23, 24, 25], dissipation is introduced by coupling to a bath of phonons that is simple enough to be traced out, thus yielding an effective description of the system of interest. Although this type of bath was first studied by putting it in contact with a single degree of freedom such as a spin in the spin-boson model [26, 27], several many-body 1D systems were later proposed, ranging from spin chains [28, 29, 30, 31, 32, 33] and 1D bosons [34], to glassy systems [35, 36], and with various system-bath couplings [37, 38]. In this work, we concentrate on the well-known XXZ spin chain at zero magnetization and temperature. Each spin of the chain is linked to a phonon bath à la Caldeira and Leggett, which can be seen as a many-body extension of the spin-boson model. The baths are furthermore spatially correlated by bath-bath nearest neighbor interactions. To the best of the author’s knowledge, it is the first time that such correlated dissipation is studied, previous studies having concentrated on local independent baths [31, 32, 33] or a single global bath [39]. For a spin chain at zero magnetization and with uncorrelated local baths, it has been proposed in [33] that the system presents a BKT-like transition which is driven by dissipation for subohmic baths. In this article, we broaden the scope of this study by considering correlated baths. In order to fully understand the nature of the transition in our system, we follow an approach based on a mapping to a generalized Coulomb gas. Within this gas picture, we show that for superohmic baths, the transition is the exact BKT transition while, for subohmic baths, dissipation induces a BKT-like transition. It is worth mentioning that dissipative systems such as a single [40] or an array [41, 42] of resistively shunted Josephson junctions have already been studied using approximate mappings to 1D or 2D gases with logarithmic interactions. The gaseous picture also arises naturally in the context of quantum Monte-Carlo studies of dissipative spin chains [27] where the particles are the kinks and anti-kinks of world line configurations of spins. However, our work differs from these in two ways. First, our mapping does not require any approximation and is thus mathematically exact and, second, our baths are spatially correlated and thus provide a larger variety of dissipation-induced particle-particle interactions in the gas picture.
The manuscript is organized as follows. Section 2 gives a concise summary of the main results which are derived in the following sections. Section 3 presents the microscopic model and shows how its low-energy behavior can be described by a dissipative sine-Gordon field theory. In section 4, this field theory is mapped onto a generalized 2D Coulomb gas and its basic features are described. The core of the article, identifying the BKT-like transition, is done in section 5. Finally, a brief discussion of the results and concluding remarks are made in section 6. Some technical details have been relegated in appendices A.1 to E. Throughout this article, we set $\hbar=k_{B}=1$.
@@ 3 Model
where $s>0$ is the bath exponent, $\alpha$ measures the strength of the coupling to the bath, and $1/\tau_{c}$ is a frequency cutoff above which $J(\Omega)=0$. The exact numerical prefactor in Eq. (5) is unimportant and can be found in appendix A.1. This form of $J(\Omega)$ accounts for many types of environments, ranging from acoustic-phonon baths to Kondo physics [26]. Following the classification from Caldeira and Leggett, we call $s=1$ an ohmic bath, while $10<s<$ 1 is a subohmic bath and $s>1$ is a superohmic bath. The rest of this section will be dedicated to deriving an effective action for the system. The procedure is similar to that outlined in [33, 31, 32] and relies on two key steps: first bosonizing the spin chain, and then integrating out the bath degrees of freedom.
@@ 3.2 Integrating out the bath degrees of freedom
where $K_{0}$ is a modified Bessel function of the second kind, and we have used the definition of the spectral function given in Eq. (4). Upon using the low-energy behavior of $J(\Omega)$ given in Eq. (5), one can show that $\mathcal{K}(x,\tau)\simeq\alpha\tau_{c}^{s-1}\left(\tau^{2}+x^{2}/v^{2}\right)^{-1-s/2}/u$ for $\sqrt{\tau^{2}+x^{2}/v^{2}}\gg\tau_{c}$ (see appendix A.1). Putting everything together, one arrives at the following effective interaction
@@
where $\mathcal{D}(x,\tau)=\left[(\tau/\tau_{c})^{2}+(u/v)^{2}(x/a)^{2}\right]^{-1-s/2}$ is the dimensionless dissipative kernel. This effective action is that of a Luttinger liquid (LL) with two types of interactions: a local interaction describing the spin chain’s internal dynamics and controlled by $g$, and a long-range bath-induced interaction controlled by $\alpha$. This dissipative interaction can be thought of as the bath storing information about a spin at position $x$ and time $\tau$, propagating it over a distance $x-x^{\prime}$ during a period $\tau-\tau^{\prime}$, and giving it back at time $\tau^{\prime}$ to the spin at position $x^{\prime}$. Notice that this very-strong non-Markovian effect was obtained through the exact integration of the bath and is a feature which cannot be reproduced by usual open quantum system techniques such as the Lindblad master equation. As mentioned previously, the effective action written in Eq. (3.2) has already been studied in the $v=0$ case [33] which corresponds to the limit of uncorrelated baths. In this case, the last term in Eq. (3.2) gets modified by changing the kernel into $\mathcal{D}(x,\tau)\propto\delta(x)/|\tau|^{1+s}$ (see Appendix. A.2 for the derivation). On the other hand, the case of $v\gg 1$ is new and corresponds to the limit of Markovian, but spatially correlated, baths. In this limit, the bath-induced kernel becomes $\mathcal{D}(x,\tau)\propto\delta(\tau)/|x|^{1+s}$. These two limits for $v$ formally yield identical models up to an exchange of space and time (and a redefinition of $\alpha$ to absorb extra factors of $v/u$). Another remarkable model is achieved when $v=u$ which corresponds to an isotropic system enjoying an $O(2)$ symmetry in $(x,u\tau)$ space. For the majority of this paper we will mainly deal with finite $v$ and postpone the discussion of the limiting cases ($v$ small or large) to section 6.
@@ 6 Discussion and conclusion
The procedure followed in this article can be straightforwardly extended to other one-dimensional quantum systems described by a sine-Gordon-like effective action. Studies of an XXZ spin chain, at zero [33] and finite magnetization [31, 32], subject to local baths (i.e. $v=0$ in our model) can be readily understood in the gas picture. Adapting our analysis, one would find a dissipative kernel $\mathcal{K}(x,\tau)\propto\delta(x)/|\tau|^{1+s}$ (see Appendix A.2). This corresponds to bath-induced particle-pairs with coordinates $(x,u\tau)$ and $(x,u\tau^{\prime})$ (notice the identical $x$ coordinates). While the presence of particle-pairs of fixed orientation might seem to be an artifact specific to this model, it is actually a feature shared by all field theories endowed with the statistical tilt symmetry (STS) or space-gauged shift symmetry, i.e. the invariance of the interacting part of the action $S-S_{\rm LL}$ under $\phi(x,\tau)\to\phi(x,\tau)+f(x)$. In the gas picture, the STS indeed requires that the total charge along any $x={\rm cst}$ slice of the gas must be zero, thus strongly restricting the possible relative orientations of particles. This result is easily shown by considering the Boltzmann weight $\mathcal{W}\propto\langle\prod_{j}\exp[i4q_{j}\phi(r_{j})]\rangle_{\rm LL}$ associated with a configuration of particles of charges $q_{j}$ at positions $r_{j}$. Using the STS, one gets that $\mathcal{W}\propto\exp[i4\sum_{j}q_{j}f(x_{j})]\langle\prod_{j}\exp[i4q_{j}\phi(r_{j})]\rangle_{\rm LL}$ for all functions $f(x)$. This implies that allowed particle configurations must be neutral along any $x={\rm cst}$ slice of the 2D space. The STS appears generically in the replicated action of one-dimensional systems with quenched disorder, as in the well-known Giamarchi-Schulz model [16, 19].
@@ Acknowledgements
Appendix A Bath kernel
A.1 General derivation
@@
where the numerical prefactor is $N^{-1}=\frac{u}{v\pi^{2}}\int_{0}^{\infty}{\rm d}y\,y^{1+s}K_{0}(y)$. Note that $N$ is positive since $K_{0}(y)>0$ for $y\in\mathbbm{R}_{+}$, and well-defined for all $s\geq 0$ as $K_{0}(y)\overset{y\gg 1}{\sim}\sqrt{\frac{\pi}{2y}}e^{-y}$. Thus,
$\mathcal{K}(x,\tau)=$ $\frac{2}{\pi}\int{\rm d}\Omega\,J(\Omega)\Omega\frac{1}{2\pi v}K_{0}\left(\Omega\sqrt{\tau^{2}+x^{2}/v^{2}}\right)$
$=$ $\frac{N\alpha\tau_{c}^{s-1}}{\pi^{2}v}\int_{0}^{1/\tau_{c}}{\rm d}\Omega\,\Omega^{1+s}K_{0}\left(\Omega\sqrt{\tau^{2}+x^{2}/v^{2}}\right)$
$=$ $\frac{\alpha\tau_{c}^{s-1}}{u\sqrt{\tau^{2}+x^{2}/v^{2}}^{2+s}}\frac{\int_{0}^{\sqrt{\tau^{2}+x^{2}/v^{2}}/\tau_{c}}{\rm d}y\,y^{1+s}K_{0}(y)}{\int_{0}^{\infty}{\rm d}y\,y^{1+s}K_{0}(y)}$
@@
A.2 Uncorrelated and Markovian limits
We derive the kernels $\mathcal{K}(x,\tau)$ obtained in the limit of uncorrelated baths $v\to 0$ and in the Markovian limit $v\to\infty$. Let us first proceed with the $v\to 0$ limit. We start from the second line of Eq. (A.1),
[…]
and assume that $\tilde{\alpha}=\alpha v/u\propto\alpha N$ is kept constant when $v\to 0$ (which, from Eq. (A.2) just amounts to $J(\Omega)$ having a well-defined limit when $v\to 0$). If we let $v\to 0$ for $x\neq 0$, because of the exponential decay of the modified Bessel function $K_{0}$ for large argument, we obtain $\lim_{v\to 0}\mathcal{K}(x\neq 0,\tau)=0$, while it is obvious that for $x=0$, $\lim_{v\to 0}\mathcal{K}(0,\tau)=+\infty$. To verify that $\mathcal{K}(x,\tau)$ behaves as a Dirac $\delta(x)$ distribution, we must calculate the weight
$\int_{-\infty}^{+\infty}{\rm d}x\mathcal{K}(x,\tau)=$ $\frac{N\alpha\tau_{c}^{s-1}}{\pi^{2}}\int_{0}^{1/\tau_{c}}{\rm d}\Omega\,\Omega^{1+s}\int_{-\infty}^{+\infty}\frac{{\rm d}x}{v}K_{0}\left(\Omega\sqrt{\tau^{2}+x^{2}/v^{2}}\right)$
$=$ $\frac{1}{|\tau|^{1+s}}\frac{N\alpha\tau_{c}^{s-1}}{\pi^{2}}\int_{0}^{|\tau|/\tau_{c}}{\rm d}y\,y^{1+s}\int_{-\infty}^{+\infty}{\rm d}u\,K_{0}\left(y\sqrt{1+u^{2}}\right).$ (A.5)
When $|\tau|\gg\tau_{c}$, we can extend the $y$ integration to +$\infty$+ ∞ and use the definition of the normalization $N$ to find
$\int_{-\infty}^{+\infty}{\rm d}x\mathcal{K}(x,\tau)=$ $\frac{\tilde{\alpha}\tau_{c}^{s-1}}{|\tau|^{1+s}}\frac{\int_{0}^{+\infty}{\rm d}z\,z^{1+s}K_{0}(z)\int_{-\infty}^{+\infty}{\rm d}u\,(1+u^{2})^{-1-s/2}}{\int_{0}^{+\infty}{\rm d}yy^{1+s}K_{0}(y)}$
$=$ $\frac{\tilde{\alpha}\tau_{c}^{s-1}}{|\tau|^{1+s}}\frac{\sqrt{\pi}\Gamma(\frac{1+s}{2})}{\Gamma(1+s/2)}.$ (A.6)
Therefore, $\lim_{v\to 0}\mathcal{K}(x,\tau)\propto\delta(x)\tilde{\alpha}/|\tau|^{1+s}$ and one recovers the decay $1/|\tau|^{1+s}$ expected for independent baths.
The Markovian limit $v\to\infty$ is obtained in a similar fashion by computing the weight $\int{\rm d}\tau K(x,\tau)$. One finds that
$\int_{-\infty}^{+\infty}{\rm d}\tau\mathcal{K}(x,\tau)=$ $\frac{N\alpha\tau_{c}^{s-1}v^{s}}{\pi^{2}}\frac{1}{|x|^{1+s}}\int_{0}^{|x|/(v\tau_{c})}{\rm d}y\,y^{1+s}\int_{-\infty}^{+\infty}{\rm d}uK_{0}\left(y\sqrt{1+u^{2}}\right).$ (A.7)
The interesting scaling limit is obtained by adopting the new UV cutoff $\tau_{c}^{\prime}=\tau_{c}v/u$ and keeping constant $\tilde{\alpha}=\alpha(v/u)^{2}$ such that, for distances $|x|\gg u\tau_{c}^{\prime}$,
$\int_{-\infty}^{+\infty}{\rm d}\tau\mathcal{K}(x,\tau)=$ $\frac{\tilde{\alpha}\tau_{c}^{\prime s-1}u^{s}}{|x|^{1+s}}\frac{\sqrt{\pi}\Gamma(\frac{1+s}{2})}{\Gamma(1+s/2)}.$ (A.8)
which shows that the kernel reduces to $\lim_{v\to\infty}\mathcal{K}(x,\tau)\propto\delta(\tau)\tilde{\alpha}/|x|^{1+s}$.